# Discrete Duality Symmetry in a $3D$ Field-Theoretic Model

### v2

@@ 5 Conclusions
In our present investigation, we have taken into account the combined system of the free $3D$ Abelian 1-form and 2-form gauge theories together and shown the existence of (i) the nilpotent BRST symmetry transformations, (ii) the nilpotent co-BRST symmetry transformations, and (iii) a bosonic symmetry transformation which is constructed from the anticommutator of the BRST and co-BRST symmetry transformations. Under the BRST symmetry transformations, the Lagrangian densities of the free Abelian 1-form and 2-form gauge theories remain invariant, separately and independently. However, for the validity of the co-BRST symmetry invariance, we find that (i) the kinetic term of the Abelian 2-form field, and (ii) the FP-ghost term of the Abelian 1-form theory transform in such a manner that they produce a total spacetime derivative******In exactly similar fashion, we note that (i) the kinetic term of the Abelian 1-form gauge field, and (ii) the basic FP-ghost term of the Abelian 2-form theory, transform to a total spacetime derivative under the co-BRST symmetry transformations thereby leading to the invariance of the action integral.. The algebra obeyed by these symmetry operators [cf. Eq. (4)] is reminiscent of the algebra obeyed by the de Rham cohomological operators of differential geometry [cf. Eq. (4)].
@@
One of the highlights of our present investigation is the existence of the discrete duality symmetry transformations in our BRST-quantized theory which has its origin in the Hodge duality operator (on the 3D flat Minkowskian spacetime manifold) as far as the basic gauge fields of the Abelian 1-form and 2-form theories are concerned [cf. Eq. (16) and corresponding footnote]. The interplay between the discrete duality symmetry transformations [cf. Eqs. (3),(3)] and continuous symmetry transformations [cf. Eqs. (3),(3)] for the (co-)BRST invariant Lagrangian density (21) are responsible for the physical realization (29) of the well-known relationship (i.e. $\delta=\pm\,*\,d\,*$) between the (co-)exterior derivatives of differential geometry [14-18]. The most basic relationship (16) between the gauge fields of the Abelian 1-form and 2-form theories establishes clearly that there is a duality symmetry between the free Abelian 2-form and 1-form basic fields††††††We would like to mention that these duality symmetry transformations have not been mentioned in our very recent work [20] on the combined system of the free $3D$ Abelian 1-form and 2-form gauge theories. This is why we have not been able to obtain the analogue of (29). in three (i.e. $D=3$) dimensions of spacetime when they are taken together for specific purpose. It interesting to add that these discrete duality symmetry transformations [cf. Eqs. (3),(3)] are also responsible for establishing a direct connection between the off-shell nilpotent (co-)BRST symmetry transformation operators (cf. Appendix A for details).
@@
We would like to say a few words about the presence of the scalar and pseudo-scalar fields (i.e. $\phi,\,\widetilde{\phi}$) in our (co-)BRST invariant Lagrangian density (21). It is very interesting to observe that (i) the scalar field appears in our theory with a positive kinetic term, and (ii) the pseudo-scalar field is endowed with a negative kinetic term (cf. Sec. 1 for more details. The latter field (having the negative kinetic term) is one of a possible candidates for the “phantom” fields which have become quite popular in the context of the cyclic, bouncing and self-accelerated cosmological models of the Universe (see, e.g. [21-25] and references therein). Such massless fields (having only the negative kinetic term) are also relevant in the context of being one of the passable candidates for dark energy (see, e.g. [26,27] for details) which are supposed to be responsible for the current observations of the accelerated cosmic expansion [32-35] of the Universe.
@@
In our present investigation, we have (i) not computed the Noether conserved charges corresponding to the off-shell nilpotent (co-)BRST symmetry transformations [cf. Eqs. (3),(3)] as well as the bosonic symmetry transformations [cf. Eq. (4)], and (ii) not considered the generalization of the classical gauge-fixed Lagrangian density (19) to its quantum counterpart within the framework of BRST formalism. In his context, it is gratifying to state that, in our very recent paper [36], we have considered the quantum generations of the both the gauge-fixed Lagrangian densities (18) and (19) and obtained the two-to-one mappings between the full set of symmetry transformation operators of the BRST-quantized theory and the set of cohomological operators of differential geometry. In our future endeavors, we plan to devote time to address the above issues and compute the extended BRST algebra among the appropriate set of conserved charges of our theory and establish its deep connection with the algebra obeyed by the de Rham cohomological operators of differential geometry so that we can prove that our present odd dimensional (i.e. $3D$) field-theoretic model of the combined system of the free Abelian 1-form ad 2-form gauge theories is a perfect example for Hodge theory.
@@ Appendix A On Direct Connection: Nilpotent $s_{d}$ and $s_{b}$
Besides the algebraic relationship we have obtained between the off-shell nilpotent ($s_{d}^{2}=0,\;s_{b}^{2}=0$) (co-)BRST symmetry transformation operators $(s_{d})s_{b}$ in equation (29), there is a direct relationship between these fermionic (i.e. off-shell nilpotent) symmetry transformation operators due to the presence of the discrete duality symmetry transformation operators [cf. Eqs. (3),(3)]. To corroborate this statement, let us pick-up the nilpotent BRST symmetry transformation: $s_{b}\,B_{\mu\nu}=-\,(\partial_{\mu}C_{\nu}=\partial_{\nu}C_{\mu})$ and apply the discrete duality symmetry transformation operators (3) and (3) on it as follows:
@@
Using the discrete transformations (3) and (3) and taking into account the operation: $*\,s_{b}=s_{d}$, we obtain the following explicit relationship
$s_{d}\Big{(}\pm\,i\,\varepsilon_{\mu\nu\sigma}\,A^{\sigma}\Big{)}=-\,\Big{[}\partial_{\mu}\big{(}\pm\,i\,\bar{C}_{\mu}\big{)}-\partial_{\nu}\big{(}\pm\,i\,\bar{C}_{\mu}\big{)}\Big{]},$ (A.2)
@@
Thus, starting from the BRST symmetry transformation for an antisymmetry tensor gauge field $B_{\mu\nu}$, we have obtained the co-BRST symmetry transformation for the Abelian 1-form vector field $A_{\mu}$ correctly due to the direct application of the discrete duality symmetry transformation operators [cf. Eqs. (3),(3)] and taking into account the discrete duality symmetry transformation for the BRST symmetry transformation operator $s_{b}$ as: $*\,s_{b}=s_{d}$ (which leads to the derivation of the co-BRST symmetry transformation for $A_{\mu}$ field).
@@
where we take: $*\,s_{d}=-\,s_{b}$ as an input. We observe the following from (A.4), namely;
@@
Thus, we conclude that the off-shell nilpotent ($s_{d}^{2}=0,\;s_{b}^{2}=0$) (co-)BRST symmetry transformation operators $(s_{d})s_{b}$ are directly connected if we use the duality property: $*\,s_{b}=s_{d},\;*\,s_{d}=-\,s_{b}$ and the discrete duality symmetry transformation operators [cf. Eqs. (3),(3)] judiciously. Ultimately, we have been able to establish that the mapping: $s_{b}\Leftrightarrow s_{d}$ can be realized (if we use directly the discrete duality symmetry transformation operators [cf. Eqs. (3),(3)] along with the inputs: $*\,s_{b}=s_{d}$,  $*\,s_{d}=-\,s_{b}$). We have been able to establish these kinds of relationships among all the (co-)BRST symmetry transformation operators $(s_{d})s_{b}$ [cf. Eqs. (3),(3)] as follows:
@@
We end this Appendix with the following remarks. First, we note that the basic discrete duality symmetry transformations [cf. Eq. (16)] exist between the Abelian 1-form field $A_{\mu}$ and 2-form field $B_{\mu\nu}$ at the level of classically gauge-fixed Lagrangian density [cf. Eq. (12)]. Second, these duality relationships are promoted at the quantum level when the Faddeev-Popov (FP) ghost terms are added to the gauge-fixed Lagrangian density where we have to take the total discrete duality symmetry transformations for the non-ghost fields [cf. Eq. (3)] as well as for the (anti-)ghost fields [cf. Eq. (3)]. Third, we have noted in this Appendix that the total set of discrete duality symmetry transformations [cf. Eqs. (3),(3)] provide a direct mapping between the off-shell nilpotent BRST and co-BRST symmetry transformation operators [cf. Eq. (A)] provided we use: $*\,s_{b}=s_{d}$,  $*\,s_{d}=-\,s_{b}$. Finally, it is very interesting to point out that the duality symmetry transformations between the BRST and co-BRST symmetry operators (i.e. $*\,s_{b}=s_{d}$,  $*\,s_{d}=-\,s_{b}$) are just like the electromagnetic duality symmetry transformations between the electric and magnetic fields of the source-free Maxwell’s theory of electromagnetism in vacuum.

### v4

@@ 5 Conclusions
In our present investigation, we have taken into account the combined system of the free $3D$ Abelian 1-form and 2-form gauge theories together and shown the existence of (i) the nilpotent BRST symmetry transformations, (ii) the nilpotent co-BRST symmetry transformations, and (iii) a bosonic symmetry transformation which is constructed from the anticommutator of the BRST and co-BRST symmetry transformations. Under the BRST symmetry transformations, the Lagrangian densities of the free Abelian 1-form and 2-form gauge theories remain invariant, separately and independently. However, for the validity of the co-BRST symmetry invariance, we find that (i) the kinetic term of the Abelian 2-form field, and (ii) the FP-ghost term of the Abelian 1-form theory transform in such a manner that they produce a total spacetime derivative††††††In exactly similar fashion, we note that (i) the kinetic term of the Abelian 1-form gauge field, and (ii) the basic FP-ghost term of the Abelian 2-form theory transform to a total spacetime derivative under the co-BRST symmetry transformations thereby leading to the invariance of the action integral.. The algebra obeyed by these symmetry operators [cf. Eq. (4)] is reminiscent of the algebra satisfied by the de Rham cohomological operators of differential geometry [cf. Eq. (4)].
@@
One of the highlights of our present investigation is the existence of the discrete duality symmetry transformations in our BRST-quantized theory which has its mathematical origin in the Hodge duality operator (on the $3D$ flat Minkowskian spacetime manifold) as far as the basic gauge fields of the Abelian 1-form and 2-form theories are concerned [cf. Eq. (16) and corresponding footnote]. The interplay between the discrete duality symmetry transformations [cf. Eqs. (3), (3)] and continuous symmetry transformations [cf. Eqs. (3), (3)] for the (co-)BRST invariant Lagrangian density (21) are responsible for the physical realization [cf. Eq. (29)] of the well-known relationship (i.e. $\delta=\pm\,*\,d\,*$) between the (co-)exterior derivatives of differential geometry [14, 15, 16, 17, 18]. The most basic relationship (16) between the gauge fields of the Abelian 1-form and 2-form theories establishes clearly that there is a duality symmetry between the free Abelian 2-form and 1-form basic fields‡‡‡‡‡‡We would like to mention that these duality symmetry transformations have not been mentioned in our very recent work [20] on the combined system of the free $3D$ Abelian 1-form and 2-form gauge theories. This is why we have not been able to obtain the analogue of (29). in three (i.e. $D=3$) dimensions of spacetime when they are taken together for specific purpose. It is interesting to add that these discrete duality symmetry transformations [cf. Eqs. (3), (3)] are also responsible for establishing a direct connection between the off-shell nilpotent (co-)BRST transformation operators (cf. A for details).
@@
We would like to say a few words about the presence of the scalar and pseudo-scalar fields (i.e. $\phi,\,\widetilde{\phi}$) in our (co-)BRST invariant Lagrangian density (21). It is very interesting to observe that (i) the scalar field appears in our theory with a positive kinetic term, and (ii) the pseudo-scalar field is endowed with a negative kinetic term (cf. Sec. 1 for more details). The latter field (having the negative kinetic term) is one of the possible candidates for the “phantom” fields which have become quite popular in the context of the cyclic, bouncing and self-accelerated cosmological models of the Universe (see, e.g. [21, 22, 23, 24, 25] and references therein). Such massless fields (having only the negative kinetic term) are also relevant in the context of being one of the passable candidates for dark energy (see, e.g. [26, 27] for details) which are supposed to be responsible for the current observations of the accelerated cosmic expansion [32, 33, 34, 35] of the Universe.
As far as our present $3D$ field-theoretic system is concerned, the pseudo-scalar field appears in the generalization of the kinetic term $\frac{1}{4}\,(F^{\mu\nu})^{2}=-\,\frac{1}{2}\,(\varepsilon_{\mu\nu\sigma}\,\partial^{\nu}A^{\sigma})^{2}$] of the Abelian 1-form theory when it is expressed in terms of the Hodge dual of the Abelian 2-form: $F^{(2)}=*\,\big{[}\frac{1}{2!}\,F_{\mu\nu}\,(dx^{\mu}\wedge dx^{\nu})\big{]}\equiv(\varepsilon_{\mu\nu\sigma}\,\partial^{\nu}A^{\sigma})\,dx^{\mu}$ which turns out to be a pseudo 1-form because the quantity ($\varepsilon_{\mu\nu\sigma}\,\partial^{\nu}A^{\sigma}$) is an axial-vector. Here we have the freedom to add/subtract [cf. Eqs. (18), (19)] a pseudo 1-form $\widetilde{\phi}^{(1)}=(\partial_{\mu}\widetilde{\phi})\,dx^{\mu}$ that is constructed from the derivative on a pseudo-scalar field $\widetilde{\phi}$. This massless pseudo-scalar field (with the negative kinetic term) is very important in our theory because it plays a crucial role in the existence of (i) the discrete duality symmetry transformations [cf. Eq. (3)], and (ii) the non-trivial dual-BRST (i.e. co-BRST) symmetry transformations [cf. Eq. (3)]. A close look at the equations (18), (19), (3), (3) and A demonstrates that we have the nilpotent co-BRST symmetry transformations in our combined field-theoretic system of the free Abelian 2-form and 1-form gauge theory because of the presence of the pseudo-scalar field. In contrast, the pure scalar field remains inert as far as the co-BRST symmetry transformations are concerned [cf. Eq. (3)]. We would like to re-emphasize that the existence of (i) the nilpotent co-BRST symmetry transformations, and (ii) the pseudo-scalar field, is crucial for our $3D$ theory to be a field-theoretic example for Hodge theory within the framework of BRST formalism.
@@
In our present investigation, we have (i) not computed the Noether conserved charges corresponding to the off-shell nilpotent (co-)BRST symmetry transformations [cf. Eqs. (3), (3)] as well as the bosonic symmetry transformations [cf. Eq. (4)], and (ii) not considered the generalization of the classical gauge-fixed Lagrangian density (19) to its quantum counterpart within the framework of BRST formalism. In this context, it is gratifying to state that, in our very recent paper [36], we have considered the quantum generalizations of both the gauge-fixed Lagrangian densities (18) and (19) and obtained the two-to-one mappings between the full set of symmetry transformation operators of the BRST-quantized $3D$ theory and the set of cohomological operators of differential geometry. In our future endeavors, we plan to devote time to address the above issues and compute the extended BRST algebra among the appropriate set of conserved charges of our $3D$ theory and establish its deep connection with the algebra obeyed by the de Rham cohomological operators of differential geometry [14, 15, 16, 17, 18] so that we can prove that our present odd dimensional (i.e. $3D$) field-theoretic model is a perfect example for Hodge theory. This proof will be different from our earlier works [6, 7, 8, 9] that are in the even dimensions of spacetime.
@@ Acknowledgment
Data Availability
No data were used to support this study.
Conflicts of Interest
The authors declare that there are no conflicts of interest.
@@ Appendix A On Direct Connection: Nilpotent $s_{d}$ and $s_{b}$
Besides the algebraic relationship we have obtained between the off-shell nilpotent ($s_{d}^{2}=0,\;s_{b}^{2}=0$) (co-)BRST symmetry transformation operators $(s_{d})s_{b}$ in equation (29), there is a direct relationship between these fermionic (i.e. off-shell nilpotent) symmetry transformation operators due to the presence of the discrete duality symmetry transformation operators [cf. Eqs. (3), (3)]. To corroborate this statement, let us pick-up the nilpotent BRST symmetry transformation: $s_{b}\,B_{\mu\nu}=-\,(\partial_{\mu}C_{\nu}-\partial_{\nu}C_{\mu})$ and apply the discrete duality symmetry transformation operators (3) and (3) on it as follows:
@@
Using the discrete transformations (3) and (3) and taking into account the operation: $s_{b}=s_{d}$, we obtain the following explicit relationship
$s_{d}\Big{(}\pm\,i\,\varepsilon_{\mu\nu\sigma}\,A^{\sigma}\Big{)}=-\,\Big{[}\partial_{\mu}\big{(}\pm\,i\,\bar{C}_{\nu}\big{)}-\partial_{\nu}\big{(}\pm\,i\,\bar{C}_{\mu}\big{)}\Big{]},$ (A.2)
@@
Thus, starting from the BRST symmetry transformation for an antisymmetry tensor gauge field $B_{\mu\nu}$, we have obtained the co-BRST symmetry transformation for the Abelian 1-form vector field $A_{\mu}$ correctly due to the direct application of the discrete duality symmetry transformation operators [cf. Eqs. (3), (3)] and taking into account the discrete duality symmetry transformation for the BRST symmetry transformation operator $s_{b}$ as: $s_{b}=s_{d}$ (which leads to the derivation of the co-BRST symmetry transformation for $A_{\mu}$ field).
@@
where we take: $s_{d}=-\,s_{b}$ as an input. We observe the following from (A.4), namely;
@@
Thus, we conclude that the off-shell nilpotent ($s_{d}^{2}=0,\;s_{b}^{2}=0$) (co-)BRST symmetry transformation operators $(s_{d})s_{b}$ are directly connected if we use the duality property: $s_{b}=s_{d},\;*\,s_{d}=-\,s_{b}$ and the discrete duality symmetry transformation operators [cf. Eqs. (3), (3)] judiciously. Ultimately, we have been able to establish that the mapping: $s_{b}\Leftrightarrow s_{d}$ can be realized (if we use directly the discrete duality symmetry transformation operators [cf. Eqs. (3), (3)] along with the inputs: $s_{b}=s_{d}$,  $s_{d}=-\,s_{b}$). We have been able to establish these kinds of relationships among all the (co-)BRST symmetry transformation operators $(s_{d})s_{b}$ [cf. Eqs. (3), (3)] as follows:
@@
We end this Appendix with the following remarks. First, we note that the basic discrete duality symmetry transformations [cf. Eq. (16)] exist between the Abelian 1-form field $A_{\mu}$ and 2-form field $B_{\mu\nu}$ at the level of classically gauge-fixed Lagrangian density [cf. Eq. (12)]. Second, these duality relationships are promoted at the quantum level when the Faddeev-Popov (FP) ghost terms are added to the gauge-fixed Lagrangian density where we have to take the total discrete duality symmetry transformations for the non-ghost fields [cf. Eq. (3)] as well as for the (anti-)ghost fields [cf. Eq. (3)]. Third, we have noted in this Appendix that the total set of discrete duality symmetry transformations [cf. Eqs. (3),(3)] provide a direct mapping between the off-shell nilpotent BRST and co-BRST symmetry transformation operators [cf. Eq. (A)] provided we use: $s_{b}=s_{d}$,  $s_{d}=-\,s_{b}$. Finally, it is very interesting to point out that the duality symmetry transformations between the BRST and co-BRST symmetry operators (i.e. $s_{b}=s_{d}$,  $s_{d}=-\,s_{b}$) are just like the electromagnetic duality symmetry transformations between the electric and magnetic fields of the source-free Maxwell’s theory of electromagnetism in vacuum.